A stone is thrown by 30°with velocity of projection 40m/s. Find {1} time of flight {2} maximum height {3} horizontal range

Grade:11

1 Answers

Aashu Mittal
6 years ago
Time of flight=2usin30°/g. T=2*40*sin30°/g=80/20=4sec. Ans...(Taking g=10m/sec^2)Maximum Height=u^2(sin30°)^2/2gH=1600/80=20m Ans....Horizontal range R= u^2sin2.30°/gH=1600sin60°/g=1600√3/20=80√3m Ans...